# MAIN: Mutual Alignment Is Necessary for instruction tuning

### v1

@@ 4.2 Ablation Study
We perform further ablation studies to analyze the effect of filtering stragety and dynamic weighting.
@@
Effective filtering plays a crucial role in improving alignment quality by removing noisy or misaligned instruction-response pairs. Table 2 presents a results comparing models trained with no filtering, score-based filtering, and our mutual-filter approach. Our mutual-filter method achieves the highest win rate, surpassing both no filtering and score-based filtering. It consistently improves instruction-following accuracy, demonstrating that our simple mutual-filter approach, without additional model-based scoring, effectively selects high-quality instruction-response pairs for fine-tuning.
@@
Our mutual filter, by leveraging mutual-alignment models, directly favors instruction-response pairs with strong semantic coherence. By eliminating misaligned samples without requiring additional supervision, it ensures a more effective training dataset, resulting in improved instruction-following and generalization capabilities.
@@
Ours + score-Based filtering 55.26 17.63 20.21 29.11 33.72
Ours + mutual filter 58.20 20.22 23.36 31.17 35.37
Table 2: Performance evaluation of LLaMA-2-7B fine-tuned on the Falcon-RefinedWeb dataset (20K samples) under three filtering conditions: no filtering, score-based filtering, and our proposed mutual filter. All conditions operate on instruction-response pairs generated by the same reverse model.
@@
Dynamic Weighting.
As shown in Figure 4, increasing $\alpha$ generally improves instruction-following ability and output quality by emphasizing well-aligned pairs. However, excessively high $\alpha$ makes the model overly reliant on generated instruction-response pairs, leading to unstable training and degraded performance.
To mitigate this, our dynamic weighting strategy adaptively balances aligned and seed data, preventing instability while maintaining strong alignment. The results show that this approach significantly improves output quality and instruction-following.
@@ 4.3 Case Study
In both cases, the reverse model’s MAIN Instructions are more focused and specific, resulting in responses that are better aligned with the intended context. In contrast, the baseline instructions are more general. These examples demonstrate that our method generates instructions that are more closely aligned with the responses, improving instruction-response coherence and enhancing the overall quality of the generated content.
@@ 5.2 Data Generation
Datasets curated by domain experts, like OpenAssistant Conversations (Köpf et al., 2024) and Databricks Dolly-15k (Conover et al., 2023), are high quality but costly. Crowdsourced platforms like ShareGPT (Chiang et al., 2023) also contribute valuable data, especially user-uploaded conversations.
@@ 6 Conclusion
In this paper, we redefine the alignment between instructions and responses, emphasizing its crucial role in optimizing instruction tuning for LLMs. By introducing the Mutual Alignment Framework, we present an innovative approach that iteratively optimizes both instructions and responses to enhance their alignment. Additionally, we propose mutual filter, a simple yet effective method for selecting instruction-response pairs with superior alignment. Our experimental results demonstrate that models fine-tuned within this framework outperform baselines across multiple evaluation benchmarks, including AlpacaEval, IFEval, and OpenLLM. These findings highlight the importance of mutual alignment in instruction tuning and how it can optimize fine-tuning data to enhance model performance. In conclusion, this work offers valuable insights for refining instruction-response pairs, enabling more efficient and scalable instruction tuning in future LLM developments.
@@
The mutual alignment approach requires more computational resources and time compared to traditional supervised fine-tuning methods. Specifically, the iterative optimization process can be computationally intensive, particularly for very large models. Due to these resource demands, we have evaluated the effectiveness of our method only on the LLaMA-2-7B and Mistral-7B models. In future work, we will explore the scalability of our approach across larger models and a wider range of datasets to assess its generalizability and applicability. Additionally, efforts will be made to develop more efficient algorithms for optimizing instruction-response alignment and adapting the framework to various tasks and domains.
@@ Appendix A Training Details
Appendix B Ablation of Iteration.

### v2

@@ 4.2 Ablation Study
We conduct additional ablation study to analyze the impact of our filtering strategy. Further ablation experiments are provided in Appendix C.
@@
Effective filtering is critical for improving alignment quality by removing noisy or misaligned instruction–response pairs. Table 2 compares models trained with no filtering, score-based filtering, and our proposed mutual-filter method. Among these, our mutual-filter achieves the highest win rate and delivers the most consistent gains in instruction-following accuracy.
@@
Our mutual filter, by leveraging mutual-alignment models, directly favors instruction-response pairs with strong coherence. By eliminating misaligned samples without requiring additional supervision, it ensures a more effective training dataset, resulting in improved instruction-following and generalization capabilities.
@@
Ours w/ score-based filtering 55.26 17.63 20.21 29.11 33.72
Ours w/ mutual filter 58.20 20.22 23.36 31.17 35.37
Table 2: Performance evaluation of LLaMA-2-7B fine-tuned on the Falcon-RefinedWeb dataset (20K samples) under three filtering conditions. All conditions operate on instruction-response pairs generated by the same reverse model.
@@
4.3 Data Alignment Analysis
We assessed the alignment quality of instructions generated by our MAIN method compared to baselines using blind pairwise evaluations conducted by GPT-4. Specifically, we randomly selected 1000 responses from the Falcon RefinedWeb dataset. For each response, GPT-4 evaluated two candidate instructions—one from MAIN and one from a baseline—in random order to avoid positional bias. We then calculated Win, Tie, and Loss rates based on GPT-4’s judgments: Win indicates GPT-4 preferred MAIN, Tie indicates no clear preference, and Loss indicates GPT-4 preferred the baseline. Using the Qwen2.5-14B as the base model, MAIN consistently outperformed all baselines, with win rates ranging from 61.7% to 81.6%, demonstrating superior instruction alignment capability (see Table 3). The evaluation prompt used by GPT-4 is detailed in Appendix F.
Baseline Win Rate Tie Rate Loss Rate $\Delta$
Humpback 69.3% 18.6% 12.1% +56.2%
Longform 81.6% 8.6% 8.8% +72.8%
Dog Instruct 61.7% 15.1% 23.2% +38.5%
Better Alignment 64.7% 12.4% 22.9% +41.8%
Table 3: GPT-4 pairwise evaluation comparing MAIN with four baseline methods on instruction alignment quality. Each comparison is based on 1000 examples sampled from the Falcon RefinedWeb dataset. $\Delta$ denotes the win rate margin of MAIN over each baseline.
@@ 4.4 Case Study
In both cases, MAIN Instructions are more focused and specific, resulting in responses that are better aligned with the intended context. In contrast, the baseline instructions are more general. These examples demonstrate that our method generates instructions that are more closely aligned with the responses.
@@ 5.2 Data Generation
Datasets curated by domain experts, like OpenAssistant Conversations (Köpf et al., 2024) and Databricks Dolly-15k (Conover et al., 2023), are high quality but costly. Crowd-sourced platforms like ShareGPT (Chiang et al., 2023) also contribute valuable data, especially user-uploaded conversations.
@@ 6 Conclusion
In this paper, we highlight the critical role of instruction-response alignment in instruction tuning for LLMs. We introduce the Mutual Alignment Framework, which iteratively optimizes both instructions and responses to improve their coherence, along with a mutual filtering strategy to select high-quality pairs. Experiments across multiple benchmarks show that our framework enables state-of-the-art performance. These findings highlight the importance of mutual alignment in instruction tuning and offer a new perspective for refining instruction-response pairs, paving the way for more effective and principled instruction tuning in future LLM development.
@@
Our experiments cover a broad range of model families and parameter scales, demonstrating the robustness of the proposed framework across architectures and sizes. However, we have not yet evaluated MAIN on very large models due to resource limit, which may exhibit qualitatively different behaviors. This limits our ability to fully assess the potential of mutual alignment at extreme scales. Investigating its effectiveness in such settings remains an important direction for future work.
@@ Appendix A Training Details
Appendix B Generalization Results
Table 5 presents extended results that demonstrate both the scalability across varying model sizes across diverse architectures of our MAIN method. These capabilities are evidenced by MAIN’s enhanced performance compared to the baseline (Nguyen et al., 2024) on the IFEval benchmark when evaluated on models such as Qwen2.5-3B and LLaMA-2-13B.
Model Method P-S (%) I-S (%) P-L (%) I-L (%)
Qwen2.5-3B Baseline $32.13$ $35.98$ $44.02$ $47.10$
MAIN $35.78$ $39.12$ $47.42$ $51.21$
LLaMA-2-13B Baseline $16.12$ $18.72$ $28.54$ $30.20$
MAIN $18.98$ $22.63$ $30.87$ $34.85$
Table 5: Comparative instruction-following performance of MAIN and Baseline on the IFEval dataset across diverse models.
Further evaluations in Table 6 demonstrate that our MAIN framework consistently outperforms Baseline (Nguyen et al., 2024) across diverse NLP benchmarks (BLEU, ROUGE-L, SQuADv2).
Base Model Method BLEU ROUGE-L SQuADv2 (EM / F1)
Llama-2-7B Baseline $43.94$ $43.82$ 10.82 / 19.62
MAIN $47.37$ $46.39$ 12.99 / 21.30
Mistral-7B Baseline $40.75$ $39.41$ 14.53 / 21.72
MAIN $43.15$ $44.70$ 17.11 / 23.49
Table 6: Comparative performance of MAIN and Baseline on diverse NLP benchmarks.
Appendix C Ablations
Training iterations.
@@ Appendix C Ablations
Dynamic Weighting.
As shown in Figure 5, increasing $\alpha$ generally improves instruction-following ability and output preference by emphasizing well-aligned pairs. However, excessively high $\alpha$ makes the model overly reliant on generated instruction-response pairs, leading to unstable training and degraded performance.
To mitigate this, our dynamic weighting strategy adaptively balances aligned and seed data, preventing instability while maintaining strong alignment. The results show that this approach significantly improves output preference and instruction-following.
Appendix D Detailed Benchmark Settings
Dataset Metric Number of Shots
ARC_C Acc_norm 25
TruthfulQA Mc2 Zero-shot
Winogrande Acc 5
HellaSwag Acc 10
MMLU Acc_norm 5
AlpacaEval Win_rate Zero-shot
IFEval P-S, etc. Zero-shot
Table 8: Evaluation settings and key metrics for benchmark datasets under few-shot and zero-shot conditions.
We have evaluated our method under both few-shot and zero-shot conditions. Specifically, tasks such as ARC_C, HellaSwag, Winogrande, and MMLU were tested with a few-shot setup, whereas AlpacaEval, TruthfulQA, and IFEval benchmarks were evaluated under zero-shot conditions. Detailed settings and results are provided in Table 8
Appendix E Computational Cost Analysis
This section details the computational demands of our proposed method MAIN. While MAIN involves an iterative training process, the overall compute cost is carefully managed to remain modest. In each iteration, only one model is trained for a single epoch on the seed data, while the other performs inference—a lightweight operation. Typically, three iterations suffice for convergence, resulting in six training epochs across the forward and reverse models. In comparison, baseline methods (Li et al., 2023a; Nguyen et al., 2024) also train a reverse model and a ranking model for a similar number of epochs. We summarize estimated GPU hours in Table 9.
Method Models Trained Inference Needed GPU Hours
Baseline Ranking model + Reverse model No 5.0
MAIN (Ours) Forward model + Reverse model Yes 5.3
Table 9: Estimated computational cost comparison for MAIN against baselines.
Appendix F GPT-4 Evaluation Prompt
The following table presents the exact prompt used to instruct GPT-4 during blind pairwise comparisons:
GPT-4 Pairwise Evaluation Prompt
Please act as an expert evaluator of instruction-response alignment. You are given a response and two candidate instructions: Instruction A and Instruction B. Your task is to decide which instruction is better aligned with the response. Evaluate based on the following aspects: - Alignment between instruction and response
- logical consistency
- natural language fluency
Please output only one of the following:
A win — if Instruction A is clearly better aligned with the response.
B win — if Instruction B is clearly better aligned with the response.
Tie — if both instructions are equally good or equally poor.
Table 10: Prompt template used in GPT-4 pairwise evaluation of instruction-response alignment.